OpenStudy (anonymous):

Let f(x) = 3x - 6 and g(x) = x - 2. Find f/g and state its domain.

TheSmartOne (thesmartone):

Factor the 3 out of 3x-6 What do you get? :)

TheSmartOne (thesmartone):

that's the second part :)

TheSmartOne (thesmartone):

first factor out 3 from 3x-6 :)

OpenStudy (anonymous):

so thats 3(x-2) ?

TheSmartOne (thesmartone):

correct so what is \(\sf\Large\frac{3(x-2)}{x-2} = ?\)

OpenStudy (anonymous):

3, right?

TheSmartOne (thesmartone):

correct.

TheSmartOne (thesmartone):

so f(x)/g(x) = 3, where \(\sf x \neq 2\)

TheSmartOne (thesmartone):

so remember the domain is all real numbers unless we have a denominator or a square root

OpenStudy (anonymous):

OOOOOH, THANK YOU !! Especially for walking through instead of just giving the answer helped a lot :)

TheSmartOne (thesmartone):

You got the domain too, right? :)

TheSmartOne (thesmartone):

it's all real numbers except for x cannot equal to 0. even though we cross it out when we simplify, we still have to remember about it :0

TheSmartOne (thesmartone):

:) *

OpenStudy (anonymous):

yeah, thank you so much :)

TheSmartOne (thesmartone):

Anytime! :)